how do I do:
x=-3y+10
-x+y=-10
using the substitution method

Answer by wuwei96815(245)

Substitute the value of x from the first equation into the second equation:

Value of x from the first equation is x=-3y + 10

-x + y = -10 becomes -(-3y+10) + y = -10

Solve for y:
-(-3y+10) + y = -10
3y -10 + y = -10
4y = 0
y = 0

Go back and solve for x using either equation:

x = -3y + 10
x = -3(0) + 10
x = 10
So the solutions are y = 0 and x = 10.

Did I get it right?